# GoldCoin: Grounding Large Language Models in Privacy Laws via Contextual Integrity Theory

### v1

@@ B.1 Brief Introduction
The Health Insurance Portability and Accountability Act (HIPAA) of 1996 Title II sets national standards for protecting personal health information (PHI). Defined as PHI, this includes any individually identifiable health information managed by entities such as health plans, health care clearinghouses, and health care providers who transmit health information electronically. The HIPAA Privacy Rule, detailed in 45 CFR Parts 160, 162, and 164, provides federal protections for PHI, limits its disclosure without consent, and gives patients rights regarding their health information, such as accessing and amending their records. The Privacy Rule is codified at 45 CFR Part 160555https://www.ecfr.gov/current/title-45/subtitle-A/subchapter-C/part-160?toc=1 and Subparts A and E of Part 164666https://www.ecfr.gov/current/title-45/subtitle-A/subchapter-C/part-164.
@@ B.3 Details of Norm Classification
In this work, we mainly focus on the 45 CFR Part 164, which governs security and privacy concerns in the healthcare sector. Following the classification in Section 3.1, we analyze the types within each norm, the statistical results are presented in Table 6. We have identified 269 of 691 HIPAA norms that permit certain information transmission and 40 of 691 norms that prohibit specific transmission.

### v2

@@ B.1 Brief Introduction
The Health Insurance Portability and Accountability Act (HIPAA) of 1996 Title II sets national standards for protecting personal health information (PHI). Defined as PHI, this includes any individually identifiable health information managed by entities such as health plans, health care clearinghouses, and health care providers who transmit health information electronically. The HIPAA Privacy Rule, detailed in 45 CFR Parts 160, 162, and 164, provides federal protections for PHI, limits its disclosure without consent, and gives patients rights regarding their health information, such as accessing and amending their records. The Privacy Rule is codified at 45 CFR Part 160666https://www.ecfr.gov/current/title-45/subtitle-A/subchapter-C/part-160?toc=1 and Subparts A and E of Part 164777https://www.ecfr.gov/current/title-45/subtitle-A/subchapter-C/part-164.
@@ B.3 Details of Norm Classification
In this work, we mainly focus on the 45 CFR Part 164, which governs security and privacy concerns in the healthcare sector. Following the classification in Section 3.1, we analyze the types within each norm, the statistical results are presented in Table 6. Our analysis identified 269 out of 691 HIPAA norms that allow certain information transmissions, while 40 out of 691 norms prohibit specific transmissions. However, the classification of norm types by LLMs is not always accurate. As demonstrated in Table 1, there are three instances where the compliance results conflict with the specified norm type. Following expert annotation, we find out that one permit norm and two forbid norms are misclassified.
@@ F.3 Baselines under Multi-step Instruction
F.4 Robustness and Sensitivity Analysis
Due to the inconsistent quantities of permit and forbid norms in HIPAA, as depicted in Table 6, there exists a category imbalance in the generated number of cases. To ascertain whether the imbalance between norm and case categories in HIPAA adversely influences the efficacy of training and testing, we oversample 269 forbid cases (to match the number of permit cases) and compare the results with the original experiment. This involved generating ten cases for each of the 40 forbid norms and randomly selecting 269 cases from this pool, striving to maintain an equitable distribution of cases per norm. The permit type cases remain unchanged as in the uploaded dataset folder. We select two models, Mistral-7B and Llama2-13B, which perform best in the main experiments. The results, shown in Table 12, indicate that GoldCoin exhibits insensitivity to sample imbalance, suggesting that the imbalance in training data has a negligible impact on final performance.
Permit (F1) Forbid (F1) All (Ma-F1)
Mistral-7B 83.95 50.00 66.98
Mistral-7B 84.34 45.83 65.09
Llama2-13B 85.21 44.44 64.83
Llama2-13B 85.88 45.45 65.67
Table 12: Comparison of Macro F1-scores for Mistral-7B and Llama2-13B, tuned with and without oversampling data in the compliance task. Models marked with $\spadesuit$ use oversampling during tuning by GoldCoin.